Oppositely charged parallel plates are separated by 5.33mm.A potential difference of 600 V exists between the plates.(a) What is the magnitude of the electric field between the plates? (b) What is the magnitude of the force on an electron between the plates? (c) how much work must be done on the electron to move it to the negative plate if it is initially positioned 2.90mm from the positive plate?
1
Expert's answer
2020-03-23T12:23:02-0400

a)


"E=\frac{U}{d}=\frac{600}{0.00533}=113\frac{kV}{m}"

b)


"F=eE=1.6\cdot10^{-19}\frac{600}{0.00533}=1.8\cdot10^{-14}N"

c)


"W=1.8\cdot10^{-14}(0.00533-0.0029)=4.4\cdot10^{-17}\ J"